# Mapping the Trust Terrain: LLMs in Software Engineering - Insights and Perspectives

### v1

@@ 6.4. Discussion
Our findings reveal that trust in LLM4SE is shaped by an interplay of model-specific and user-centric attributes, highlighting the complexity of user confidence in LLMs. Accuracy emerged as a fundamental driver of trust. Our survey study also confirms that practitioners prioritize accuracy in LLM-generate outputs, mirroring observations from broader trust literature. However, overemphasis on accuracy without mechanisms for uncertainty expression can lead to misplaced trust, a well-documented concern in AI research. While accuracy is crucial, our findings underscore the necessity of equipping users with tools to critically assess reliability.
@@
A key gap in existing research in LLM4SE is the lack of efforts to establish priority of trust antecedents. While efforts have been made to identify by model and user-specific antecedents of trust, there has been no systematic investigation into their relative importance to trust. Understanding which trust antecedents hold the highest priority for software engineers could enable more targeted improvements in LLM4SE, ensuring that development efforts are aligned with user expectations. This need for prioritization is evident in our survey study, where users consistently emphasized certain factors, such as accuracy, interpretability, controllability, and robustness over endorsement or source reputation. A lack of clarity on the relative importance of trust antecedents could lead development efforts to focus on less critical aspects while neglecting those that matter most to users.
@@
Transparency and accountability remain central concerns for trust in LLM4SE. Our survey study indicates that experienced software engineers demand clearer explanations for model decisions, a concern echoed in broader trust literature. However, the opaque nature of LLMs, exacerbated by corporate secrecy, presents a significant barrier to transparency. Broader AI research suggests incorporating uncertainty expressions or explainable AI techniques can enhance trust. The importance of interpretability was clearly visible in our survey study, where it was deemed important across all downstream SE tasks.
@@
Robustness is another critical dimension where expectations in software engineering diverge from broader AI applications. Our findings reveal that experienced software engineers emphasize the need for LLMs to maintain consistency across different programming tasks. This aligns with concerns in AI literature, where models frequently exhibit instability when faced with adversarial inputs or minor variations in phrasing. While robustness is a universal challenge for AI, software engineering demands particularly high levels of consistency due to the necessity of reproducibility in development workflows.
@@
Ethical considerations, while not the primary concern for experienced software engineers, remain integral to trust formation in LLM4SE. Concerns about social biases in LLM-generated code and potential skill degradation among developers underscore the need for ethical safeguards. The ability of LLMs to handle diverse linguistic and cultural contexts is key to ensuring trust across the global community of Software Developers. While technical reliability is a prerequisite for trust, long-term confidence in AI systems requires ongoing attention to fairness, professional impact, and user empowerment. Although software engineers may not prioritize ethical considerations, its neglect could contribute to distrust.
@@ 6.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the trustworthiness of LLMs in the context of SE tasks. An important avenue of research could be determining the relative importance of trust antecedents to prioritize trustworthy model development. Causal analysis, survey studies, controlled experiments, and A/B testing could help quantify the impact of different factors on trust. Identifying high-priority antecedents will enable targeted development of trustworthy LLMS for various SE tasks.
@@
Another key area is community-driven trust signals, future research should investigate methods for moderating and validating community-driven feedback to ensure its reliability and usefulness. Moreover, incorporating adaptive feedback loops into LLMs could allow them to refine their performance based on real-time user input, fostering greater trust among diverse developer communities.
Lastly, tailored trust models could improve usability by adjusting trust mechanisms based on user expertise and task complexity. Novice users might benefit from simplified explanations and clear guidance, while experienced developers may seek more detailed insights into model behavior, particularly for tasks that demand high accuracy and robustness. Adapting trust signals dynamically could enhance confidence across different experience levels.
@@ 7.1. Phase1: SLR Results
User trust in AI models is shaped by perceived cues and individual experiences, varying across groups (Brown et al., 2024). According to Liao et. al.  (Liao and Sundar, 2022), users may either strategically analyze LLM-generated output or rely on heuristics like intuition for trust decisions. The goal is to cultivate calibrated trust among stakeholders, avoiding both blind acceptance and unwarranted mistrust, ensuring that trust judgments are well-founded rather than based on flawed intuition.
@@
When users lack the motivation or ability to conduct in-depth evaluations, they often form trust judgments based on surface-level cues or general impressions. Heuristics may sometimes be an intentional choice, but trust is frequently influenced by subtle, hard-to-articulate factors often linked to a “sixth sense”  (Widder et al., 2021b). Several heuristic-based trust factors have been identified in the literature.
@@
The scope of the task significantly influences trust in LLM-generated suggestions. Developers were less likely to accept multi-line suggestions for smaller code changes, which are often more straightforward and may require less assistance. In contrast, larger code changes were associated with higher acceptance rates for multi-line suggestions (Brown et al., 2024). One hypothesis is that smaller changes provide developers with clearer ideas about their intended functionality, reducing the perceived need for AI assistance. Additionally, developers were significantly less likely to accept suggestions when editing test files. This may be because test code is often constrained to specific goals that AI suggestions may not align with, or because the suggestions might not adhere to local codebase practices, favoring global best practices instead (Brown et al., 2024).
@@
On an individual level, a person’s experience in SE and LLMs, along with their disposition or mindset, influence trust judgment. Developers with recent experience using a programming language generated by LLMs are more likely to trust suggestions in that language  (Brown et al., 2024). Interestingly, novice programmers tend to exhibit higher trust in the security of machine-generated code than in human-generated code  (Perry et al., 2023). Developers trained to follow specific style guidelines for readability in a programming language are more likely to accept multi-line suggestions generated by LLMs  (Brown et al., 2024). Furthermore, users’ prior positive experiences with other models can increase their trust in new models, with their general tendency to trust further shaping their judgments  (Kästner et al., 2021)
@@
Sociological aspects of the model and its building also play a role in trust judgment. People pay attention to the reputation of the company or institution that developed the model (Noller et al., 2022; Cheng et al., 2023; Johnson et al., 2023). The amount of institutional investment is also a factor that users believe indicates the model’s trustworthiness (Widder et al., 2021b). Furthermore, the population of the model’s users can lead to people’s trust, as a larger number of users offer people assurance that the model is under many levels of supervision, encouraging their belief in the model’s security provenance (Widder et al., 2021b).
@@
Several functionality-associated factors are positively valued by users. The model is preferably easy to use (Widder et al., 2021b). For example, if the model’s features are intuitive or similar to other popular models’ usage, this allows users to operate within their comfort zone and reduces cognitive burden (Widder et al., 2021b). Also, the model’s ease of integration with the user’s existing workflow and alignment with the developer’s coding style directly assist in trust-building (Johnson et al., 2023). A polished interface design also leads to positive trust-building (Johnson et al., 2023).
@@
During interactions, users often employ heuristics to evaluate the model. For example, a sense of control over the model’s outputs and autonomy in decision-making are highly valued. Users who can directly manipulate the code and make choices about the generated results are more likely to trust the model  (Johnson et al., 2023). Over time, as users accumulate experience and exposure to the model, they become better equipped to assess its outputs, further strengthening trust  (Widder et al., 2021b). The ability of the model to generate predictable and consistent results also plays a crucial role in trust. While users may be resilient to occasional errors, they tend to lose trust in models that produce unpredictable or surprising outputs  (Widder et al., 2021b).
@@ 7.2. Phase2: Survey Study Results
Trust Metric and Measurement. For all three SE tasks, we observed that output accuracy is a highly valued aspect. Accuracy is the most frequently mentioned metric by our respondents; 38% of people who answered the question mentioned accuracy in their responses. The emphasis on accuracy highlights that users expect LLMs to deliver precise and dependable results, as reflected in the participant’s quote, ”If the generated output is flawed, there would be no point in using the model”. This expectation of output correctness aligns with the critical nature of SE, where flawed outputs can lead to significant issues.
Usability cited by 23.8% of participants, is the second most mentioned metric across tasks. It is assessed by the volume of correctly functioning code generated over time with minimal modifications needed. Users gauge trustworthiness by how well the model meets requirements without needing further changes. Usability encompasses comprehensive fulfillment of user needs, including high performance and alignment with specifications. Additionally, although LLMs are capable of generating large programs over a short period, this needs to be balanced by the ”amount of time lost trying to get a good response or wasted in debugging errors that it introduces.” With increased usability, less time needs to be spent on modifying the code, thus boosting overall productivity.
Another interesting result common across tasks is how people judge their trust in the model. This judgment is often based on their likelihood of accepting the generated output without fully understanding or carefully scrutinizing it. Respondents believe that their trust is related to their motivation to review the output. As their trust in LLMs increases, they are less likely to engage in systematic processing to gauge the model’s trustworthiness, including its ability, benevolence, and process integrity. Interestingly, these were mostly echoed by novice developers.
For Code Generation, several ability-related metrics were identified. Comprehensibility refers to the model’s ability to form logical reasoning through neural networks. Understandability pertains to the ease of understanding the code, with participants noting that their trust ”is largely tied to measures of my understanding of the generated code.” Benevolence is another metric mentioned by users, referring to the notion that ”the generated code is not maliciously trying to cause harm.”
@@
For Test Case Generation, one metric mentioned is comprehensiveness—the ability of the model to generate all possible test cases. Additionally, if the model can verify the user’s doubts, it would be deemed more trustworthy. As one respondent stated, ”I would trust the test cases more if they prove that my doubts about correct syntax and calling conventions do not apply.”
Reviewing Human versus LLM-Generated Code. Trust measurement procedures commonly include manual code reviews and running tests. Our survey revealed that 80.6% of respondents often review generated code, 16.6% rarely review it, and only 2.8% never review it. Participants often compared LLMs with human developers, suggesting that models behaving similarly to human experts are considered more trustworthy.
For code generation, 40% of participants distinguish between reviewing LLM-generated code and human-generated code. LLMs often generate code with unconventional syntax, style, and functionality, requiring more time to comprehend and integrate. One participant reported that LLMs occasionally generate ”meaningless code”, whereas ”code written by humans is usually functionally done, barring minor bugs”. However, participants acknowledge that LLM-generated code in smaller chunks can be quicker to read and process.
For test case generation, 25% of participants differentiate their review process between LLM-generated code and human-generated code, with LLM-generated tests requiring stricter scrutiny.
For program repair, 25% of participants distinguish between LLM-generated code and human-generated code reviews. Participants mentioned that the generated repair ”can often give wildly incorrect approaches (both in function and style).” Additionally, the model may produce false positive output, where the justification seems reasonable, but the change does not make sense. These issues necessitate thorough scrutiny of LLM-generated repairs. Conversely, users tend to automatically attribute more credibility to humans, assuming fewer significant flaws in their approach unless the individual has a reputation for errors.
@@ 7.3. Broader Analysis of Trust Metric
Measuring trust in LLMs for SE is a complex challenge, as it encompasses both technical and human-centric dimensions. Across AI research, trust assessment varies by domain, reflecting distinct priorities and risks. In SE and machine learning engineering, trust metrics are predominantly technical, emphasizing benchmark-driven evaluations and reproducibility  (Ahuja et al., 2020). Accuracy is often measured through adversarial testing frameworks such as FFT  (Cui et al., 2024), while TRUSTGPT assesses factual correctness by analyzing toxicity, bias, and value alignment  (Huang et al., 2023). Robustness is evaluated through adversarial success rates  (Wang et al., 2022), performance degradation from synonym substitutions  (Zhu et al., 2024), and handling of noisy inputs  (Wang et al., 2024b).
Accountability metrics in SE focus on tracking model versions, training data, and decision-making processes via immutable logs  (de Laat, 2018), while external audits further enhance reliability  (de Laat, 2018). Fairness, a growing concern in AI, is assessed using adversarial testing to detect biases in outputs  (Cui et al., 2024; Liu et al., 2023), with frameworks like TRUSTGPT standardizing toxicity and bias evaluations  (Huang et al., 2023).
In contrast, HCI research prioritizes user-centric trust metrics, relying on qualitative methods such as Likert scale assessments  (Baughan et al., 2023; Jakesch et al., 2019; Wu et al., 2023) and trust recovery studies  (Baughan et al., 2023). Transparency is evaluated through structured frameworks like POLARIS, which embeds transparency checks throughout the software development cycle  (Baldassarre et al., 2024b). Additionally, methods like LIME assess the fidelity of local explanations to overall model behavior, ensuring alignment between interpretability and performance  (Liu et al., 2021).
Despite these advancements, existing frameworks rarely integrate technical and user-centered trust metrics  (Baldassarre et al., 2024b). This gap is problematic, as trust cannot be fully established if AI systems are technically robust but socially distrusted, or user-friendly but functionally unreliable. Moreover, trust measurement frameworks often overlook the temporal dynamics of trust, which evolves over time based on user experience, failure rates, and behavioral consistency  (Lee and See, 2004b; Lewis et al., 2018). In domains such as robotics and education, trust is assessed longitudinally, capturing how user confidence fluctuates with repeated interactions. A similar approach is needed for SE, where trust in LLMs is influenced by ongoing exposure and reliability trends.

### v2

@@ 6.4. Discussion
Our findings reveal that trust in LLM4SE is shaped by an interplay of model-specific and user-centric attributes, highlighting the complexity of user confidence in LLMs. Accuracy has emerged as a fundamental driver of trust. Our survey study also confirms that practitioners prioritize accuracy in LLM-generated outputs, mirroring observations from a wider trust literature. However, overemphasis on accuracy without mechanisms for uncertainty expression can lead to misplaced trust, a well-documented concern in AI research. Although accuracy is crucial, our findings underscore the need to equip users with tools to critically assess reliability.
@@
A key gap in existing research in LLM4SE is the lack of efforts to establish a priority of antecedent trust. Although efforts have been made to identify trust antecedents by model and user specificity, there has been no systematic investigation of their relative importance to trust. Understanding which trust antecedents are the highest priority for software engineers could enable more targeted improvements in LLM4SE, ensuring that development efforts are aligned with user expectations. This need for prioritization is evident in our survey study, where users consistently emphasized certain factors, such as precision, interpretability, controllability, and robustness over endorsement or source reputation. A lack of clarity on the relative importance of trust antecedents could lead development efforts to focus on less critical aspects while neglecting those that matter most to users.
@@
Transparency and accountability remain central concerns for trust in LLM4SE. Our survey study indicates that experienced software engineers demand clearer explanations for model decisions, a concern echoed in the broader trust literature. However, the opaque nature of LLMs, exacerbated by corporate secrecy, presents a significant barrier to transparency. Broader AI research suggests that incorporating uncertainty expressions or explainable AI techniques can enhance trust. The importance of interpretability was clearly visible in our survey study, where it was deemed important in all downstream SE tasks.
@@
Robustness is another critical dimension where expectations in software engineering diverge from broader AI applications. Our findings reveal that experienced software engineers emphasize the need for LLMs to maintain consistency across different programming tasks. This aligns with concerns in the AI literature, where models frequently exhibit instability when faced with adversarial inputs or minor variations in phrasing. While robustness is a universal challenge for AI, software engineering demands particularly high levels of consistency due to the necessity of reproducibility in development workflows.
@@
Ethical considerations, while not the primary concern of experienced software engineers, remain integral to trust formation in LLM4SE. Concerns about social biases in LLM-generated code and potential skill degradation among developers underscore the need for ethical safeguards. The ability of LLMs to handle diverse linguistic and cultural contexts is key to ensuring trust across the global community of Software Developers. Although technical reliability is a prerequisite for trust, long-term confidence in AI systems requires ongoing attention to fairness, professional impact, and user empowerment.
Ethical considerations, while not always the primary concern of experienced software engineers, remain integral to trust formation in LLM4SE because they directly affect the reliability, fairness, and long-term adoption of these systems. Concerns about social biases in LLM-generated code underscore the crucial need for ethical safeguards, as such biases can lead to discriminatory software that perpetuates societal inequities and erodes public trust (Ling et al., 2025; Du et al., 2025). Furthermore, as stated in our user study, the potential for developer skill degradation is a legitimate concern for professional development. When developers become too dependent on LLM, their critical thinking and foundational competencies may erode. This erosion can foster resistance and distrust within the software engineering community, ultimately affecting the perceived value and sustainability of these tools. The importance of transparency in an LLM’s ethical guidelines, especially regarding how it handles sensitive data and makes decisions, is crucial (Zerilli et al., 2022). Without clear communication of these principles, even technically proficient systems risk breeding distrust. Research shows that transparency and accountability are the core to fostering trust in AI systems, as they enable stakeholders to understand how decisions are made and provide mechanisms for oversight and redress when needed (Du et al., 2025; Zerilli et al., 2022). Therefore, aligning LLM behavior with the principles of fairness, accountability, and transparency is not merely optional; it is a fundamental prerequisite for building enduring trust.
Beyond these ethical dimensions, our study reveals further critical areas where practitioner priorities for trust antecedents diverge from or offer new perspectives compared to established literature. For example, security emerged as an paramount concern for practitioners when defining trust, despite being largely overlooked in traditional SLR definitions of trust. Survey respondents explicitly linked trust to an LLM’s ability to produce safe, nonmalicious code and prevent vulnerabilities. This highlights a crucial gap between academic conceptualizations and the practical demands of secure software development using LLMs.
@@ 6.5. Opportunities for Future Work
The results of this study present several avenues for future research aimed at enhancing the trustworthiness of LLMs in the context of SE tasks. An important avenue of research could be determining the relative importance of trust antecedents to prioritize trustworthy model development. Causal analysis, survey studies, controlled experiments, and A/B testing could help quantify the impact of different factors on trust. Identifying high-priority antecedents will enable targeted development of a trustworthy LLMS for various SE tasks.
@@
A crucial avenue for future work is the development of concrete ethical guidelines and best practices for LLM-assisted SE. Research should focus on creating implementable strategies for practitioners and tool builders to mitigate ethical risks. This includes developing tools to help identify biases and security vulnerabilities in generated code. Future research should also advocate for designing LLMs that are more transparent about their ethical limitations, thereby integrating human oversight and accountability into the development lifecycle.
Another key area is community-driven trust signals; future research should investigate methods for moderating and validating community-driven feedback to ensure its reliability and usefulness. Moreover, incorporating adaptive feedback loops into LLMs could allow them to refine their performance based on real-time user input, fostering greater trust among diverse developer communities.
Lastly, customized trust models could improve usability by adjusting trust mechanisms based on user expertise and task complexity. Novice users might benefit from simplified explanations and clear guidance, while experienced developers may seek more detailed insights into model behavior, particularly for tasks that demand high accuracy and robustness. Adapting trust signals dynamically could improve confidence across different levels of experience.
@@ 7.1. Phase1: SLR results
User trust in AI models is shaped by perceived signals and individual experiences, varying between groups (Brown et al., 2024). According to Liao et al. (Liao and Sundar, 2022), users can either strategically analyze LLM-generated output or rely on heuristics such as intuition for trust decisions. The goal is to cultivate calibrated trust among stakeholders, avoiding both blind acceptance and unwarranted mistrust, ensuring that trust judgments are well-founded rather than based on flawed intuition.
@@
When users lack the motivation or ability to perform in-depth evaluations, they often make trust judgments based on surface-level cues or general impressions. Heuristics may sometimes be an intentional choice, but trust is often influenced by subtle, difficult-to-articulate factors often linked to the’sixth sense’ (Widder et al., 2021). Several heuristic-based trust factors have been identified in the literature.
@@
The scope of the task significantly influences the trust in LLM-generated suggestions. Developers were less likely to accept multiline suggestions for smaller code changes, which are often more straightforward and may require less assistance. In contrast, larger code changes were associated with higher acceptance rates for multiline suggestions (Brown et al., 2024). One hypothesis is that smaller changes provide developers with clearer ideas about their intended functionality, reducing the perceived need for AI assistance. Furthermore, developers were significantly less likely to accept suggestions when editing test files. This may be because test code is often restricted to specific goals that AI suggestions may not align with, or because the suggestions might not adhere to local codebase practices, instead favoring global best practices (Brown et al., 2024).
@@
On an individual level, a person’s experience in SE and LLMs, along with their disposition or mindset, influences the judgement of trust. Developers with recent experience using an LLM-generated programming language are more likely to trust suggestions in that language (Brown et al., 2024). Interestingly, novice programmers tend to show higher trust in the security of machine-generated code than in human-generated code (Perry et al., 2023). Developers trained to follow specific style guidelines for readability in a programming language are more likely to accept multi-line suggestions generated by LLMs (Brown et al., 2024). Furthermore, positive user experiences with other models can increase their trust in new models, with their general tendency to trust further shaping their judgments (Kästner et al., 2021)
@@
Sociological aspects of the model and its building also play a role in trust judgment. People pay attention to the reputation of the company or institution that developed the model (Noller et al., 2022; Cheng et al., 2023; Johnson et al., 2023). The amount of institutional investment is also a factor that users believe indicates the trustworthiness of the model (Widder et al., 2021). Furthermore, the population of model users can lead to trust from people, as a larger number of users offer people assurance that the model is under many levels of supervision, encouraging their belief in the provenance of the security of the model (Widder et al., 2021).
@@
Several factors associated with functionality are positively valued by users. The model is preferably easy to use (Widder et al., 2021). For example, if the model’s features are intuitive or similar to the usage of other popular models, this allows users to operate within their comfort zone and reduces cognitive burden (Widder et al., 2021). In addition, the ease of integration of the model with the existing user workflow and alignment with the developer’s coding style directly help build trust (Johnson et al., 2023). A polished interface design also leads to positive trust building (Johnson et al., 2023).
@@
During interactions, users often employ heuristics to evaluate the model. For example, a sense of control over the model’s outputs and autonomy in decision-making are highly valued. Users who can directly manipulate the code and make choices about the generated results are more likely to trust the model (Johnson et al., 2023). Over time, as users accumulate experience and exposure to the model, they become better equipped to assess its output, further strengthening trust (Widder et al., 2021). The ability of the model to generate predictable and consistent results also plays a crucial role in trust. Although users may be resilient to occasional errors, they tend to lose trust in models that produce unpredictable or surprising results (Widder et al., 2021).
@@ 7.2. Phase2: Survey Study Results
Trust Metric and Measurement. For all three SE tasks, we observed that the precision of the results is a highly valued aspect. Accuracy is the metric most frequently mentioned by our respondents; 38% of the people who answered the question mentioned accuracy in their responses. The emphasis on accuracy highlights that users expect LLMs to deliver precise and dependable results, as reflected in the participant’s quote ”If the generated output is flawed, there would be no point in using the model”. This expectation of output correctness aligns with the critical nature of SE, where flawed output can cause significant issues.
Usability cited by 23.8% of participants is the second most mentioned metric across tasks. It is assessed by the volume of correctly functioning code generated over time with minimal modifications needed. Users gauge trustworthiness by how well the model meets requirements without needing further changes. Usability encompasses the comprehensive fulfillment of user needs, including high performance and alignment with specifications. Additionally, although LLMs are capable of generating large programs over a short period, this needs to be balanced by the ”amount of time lost trying to get a good response or wasted in debugging errors that it introduces.” With increased usability, less time needs to be spent on modifying the code, thus boosting overall productivity.
Another interesting result common across tasks is how people judge their trust in the model. This judgment is often based on the likelihood that the author accepts the generated output without fully understanding or carefully scrutinizing it. The respondents believe that their trust is related to their motivation to review the output. As their trust in LLMs increases, they are less likely to engage in systematic processing to gauge the model’s trustworthiness, including its ability, benevolence, and process integrity. Interestingly, these were mostly echoed by novice developers.
For Code Generation, several ability-related metrics were identified. Comprehensibility refers to the model’s ability to form logical reasoning through neural networks. Understandability refers to the ease of understanding the code, and participants note that their trust ”is largely tied to measures of my understanding of the generated code.” Benevolence is another metric mentioned by users, referring to the notion that ”the generated code is not maliciously trying to cause harm.”
@@
For Test Case Generation, one metric mentioned is comprehensiveness, the ability of the model to generate all possible test cases. Additionally, if the model can verify the user’s doubts, it would be deemed more trustworthy. As one respondent stated, ”I would trust the test cases more if they prove that my doubts about correct syntax and calling conventions do not apply.”
Reviewing Human versus LLM-Generated Code. Trust measurement procedures commonly include manual code reviews and running tests. Our survey revealed that 80.6% of respondents often review generated code, 16.6% rarely review it, and only 2.8% never review it. Participants often compared LLMs with human developers, suggesting that models acting similarly to human experts are considered more trustworthy.
For code generation, 40% of the participants distinguish between reviewing LLM-generated code and human-generated code. LLMs often generate code with unconventional syntax, style, and functionality, which requires more time to comprehend and integrate. One participant reported that LLMs occasionally generate ”meaningless code”, whereas ”code written by humans is usually functionally done, barring minor bugs”. However, participants acknowledge that LLM-generated code in smaller chunks can be faster to read and process.
For the generation of test cases, 25% of participants differentiate their review process between LLM-generated code and human-generated code, with LLM-generated tests requiring stricter scrutiny.
For program repair, 25% of the participants distinguish between LLM-generated code and human-generated code reviews. Participants mentioned that the repaired ”can often give wildly incorrect approaches (both in function and style).” Furthermore, the model may produce false positive output, where the justification seems reasonable, but the change does not make sense. These issues require a thorough review of LLM-generated repairs. In contrast, users tend to automatically attribute more credibility to humans, assuming fewer significant flaws in their approach, unless the individual has a reputation for errors.
@@ 7.3. Broader Analysis of Trust Metric
Measuring trust in LLMs for SE is a complex challenge, as it encompasses both technical and human-centric dimensions. In AI research, trust assessment varies by domain, reflecting different priorities and risks. In SE and machine learning engineering, trust metrics are predominantly technical, emphasizing benchmark-driven evaluations and reproducibility (Ahuja et al., 2020). Accuracy is often measured through adversarial testing frameworks such as FFT (Cui et al., 2024), while TRUSTGPT assesses factual correctness by analyzing toxicity, bias, and alignment of values (Huang et al., 2023). Robustness is evaluated through adversarial success rates (Wang et al., 2022), performance degradation from synonym substitutions (Zhu et al., 2024), and handling of noisy inputs (Wang et al., 2024b).
Accountability metrics in SE focus on tracking model versions, training data, and decision-making processes via immutable logs (de Laat, 2018), while external audits further enhance reliability (de Laat, 2018). Fairness, a growing concern in AI, is assessed using adversarial testing to detect biases in outputs (Cui et al., 2024; Liu et al., 2023), with frameworks such as TRUSTGPT standardizing toxicity and bias evaluations (Huang et al., 2023).
In contrast, HCI research prioritizes user-centric trust metrics, relying on qualitative methods such as Likert scale assessments (Baughan et al., 2023; Jakesch et al., 2019; Wu et al., 2023) and trust recovery studies (Baughan et al., 2023). Transparency is evaluated through structured frameworks such as POLARIS, which incorporates transparency checks throughout the software development cycle (Baldassarre et al., 2024a). Furthermore, methods such as LIME assess the fidelity of local explanations to the overall behavior of the model, ensuring the alignment between interpretability and performance (Liu et al., 2021).
Despite these advancements, existing frameworks rarely integrate user-centered and technical trust metrics (Baldassarre et al., 2024a). This gap is problematic, as trust cannot be fully established if AI systems are technically robust but socially distrusted, or user-friendly but functionally unreliable. In addition, trust measurement frameworks often overlook the temporal dynamics of trust, which evolves over time based on user experience, failure rates, and behavioral consistency (Lee and See, 2004; Lewis et al., 2018). In domains such as robotics and education, trust is assessed longitudinally, capturing how user confidence fluctuates with repeated interactions. A similar approach is needed for SE, where trust in LLMs is influenced by ongoing exposure and reliability trends.